OpenStudy (anonymous):

Find the vertex of the parabola whose equation is y = -2x2 + 8x - 5. (2, 3) (2, 19) (2, 27)

OpenStudy (anonymous):

vertex formula: \(\Large \color{blue}{y=a(x-h)^2+k}\) with \(\Large (h,k)\) is vertex \(\Large y=a(x-h)^2+k=a(x^2-2xh+h^2)+k\\\Large =(a)x^2+(-2ah)x+(ah^2+k)\) Here equation is \(\Large y = -2x^2 + 8x - 5\).Thus: \(\Large a=-2\\\Large -2ah=8\\\Large ah^2+k=-5\) Therefore: \(\Large h=2\ and\ k=3\) So the vertex \(\Large (h,k)=(2,3)\)